# Using Covid-19 Response Policy to Estimate Open Water Swim Drafting Effects in Triathlon

### v1

@@ Abstract
This study investigates the causal effects of open-water swim drafting by leveraging a natural experiment induced by staggered race starts during the COVID-19 pandemic. Before 2020, athletes started in groups, enabling drafting benefits, while pandemic-related restrictions significantly reduced these opportunities. Using agglomerative hierarchical clustering of swim-out times, I analyze optimal drafting positions and estimate their impact on Swim-Out performance.
Our empirical findings reveal that swim drafting benefits were statistically insignificant in 2020 but persisted post-pandemic at slightly reduced levels. I find that drafting becomes advantageous only from the third trailing position onward, with earlier positions primarily serving to minimize fatigue. To mitigate endogeneity, I employ athlete and event fixed effects. The seemingly inverse decaying nature of drafting benefits partially addresses some concerns of simultaneous reverse causality and omitted variable bias. This study provides the first large-scale causal estimate of drafting effects in real-world triathlon race settings.
@@ Introduction
In these latter formats, the only real opportunity for drafting occurs in the swim segment, where swimmers can strategically position themselves to reduce drag. Swim drag reduction through drafting has been extensively studied in medical science, soft matter physics and sports science in Bolon et al. (2020); Chatard et al. (1990); Chatard and Wilson (2003); Ohmichi et al. (1983); Millet et al. (1999); Delextrat et al. (2003); Olbrecht (2011); Toussaint et al. (1989). However, no large-scale causal estimates of its benefits in real race scenarios have been published.
The COVID-19 pandemic introduced precautionary measures in Austria that significantly altered the swim segment dynamics. Österreichischer Triathlonverband (ÖTRV) Athletes often had to start with individually staggered swim entries, and overtaking at buoy turns was sometimes restricted. This created a natural experiment where COVID-19 functioned as an exogenous short-term shock (1–3 years) to open-water swim drafting, providing an opportunity to analyze its effects in a controlled manner.
@@
Water drafting has long been a hot topic in both sports science and professional triathlon. [With certain internet posts claiming that an athlete can gain up to between $\approx$ 20%-40% pct. on swim drafting] Beyond race performance, other strategic considerations—such as energy conservation and positioning to catch the first cycling group in elite competitions where wind drafting is permitted—play a critical role. Furthermore, wetsuit technology has advanced significantly in recent decades, and there is increasing awareness among athletes of drafting, primarily for energy preservation reasons rather than purely hydrodynamic advantages. (An evidence of this can be found in Appendices Figure 12)
@@
Hydrodynamic and Physiological Benefits of Drafting. Studies have shown that drafting significantly reduces drag and physiological strain. Bolon et al. (2020) determined that optimal drafting positions—directly behind a lead swimmer or at the hip—result in drag reductions of 40% and 30%, respectively. Chatard and Wilson (2003) found that metabolic costs, including oxygen uptake, heart rate, and blood lactate levels, were significantly lower when swimmers drafted at distances of 0 to 50 cm behind another swimmer’s toes. The most advantageous drafting distances were 0 and 50 cm, reducing passive drag by 21% and 20%, respectively. Additionally, Ohmichi et al. (1983) found that swimmers drafting directly behind another competitor completed 400-m races significantly faster than those who swam at the side.
@@
Drafting and Performance Prediction. The role of passive drag (Dp) in swim performance prediction was examined by Chatard et al. (1990), who found a strong correlation between Dp and VO2 max. Performance times were primarily influenced by VO2 max ($r=0.70$ for males, $r=0.72$ for females), but including Dp significantly improved the predictive accuracy. Anthropometric factors, including height, weight, and body surface area, also contributed to drag variability. This athlete heterogeneity underpins the need for the inclusion of athlete-specific intercepts in predictive models. Millet et al. (1999) further emphasized the cumulative impact of drafting, showing that triathletes employ energy-conserving strategies in swimming to enhance subsequent cycling and running performances.
@@
Tactical Implications for Triathletes. Drafting has been shown to enhance subsequent cycling performance. Delextrat et al. (2003) reported a 4.8% improvement in cycling efficiency when a swim was performed in a drafting position compared to an isolated swim. Millet et al. (1999) highlighted the cumulative effects of drafting in triathlon, where athletes strategically conserve energy in the swim segment to optimize performance in the cycle-run stages. Additionally, Olbrecht (2011) underscored the importance of balancing conditioning and technique, arguing that poor swimming technique results in excessive energy expenditure, ultimately affecting overall race outcomes.
@@
Impact of Equipment and External Conditions. The effect of wetsuits on swim drafting was investigated by Toussaint et al. (1989), who observed a 14% reduction in drag at typical triathlon swim speeds. This reduction, largely attributed to increased buoyancy and decreased frontal resistance, resulted in higher swim velocities. Moreover, Leitner (2021) conducted an athlete survey revealing preferences regarding mass starts versus staggered starts, highlighting the impact of the race format on drafting opportunities. The COVID-19 pandemic introduced natural experiments with staggered starts, as documented by Österreichischer Triathlonverband (ÖTRV), which temporarily altered swim drafting dynamics.
@@
Summary of Existing Literature and Conclusion. The body of research highlights drafting as a critical factor in competitive swimming and triathlon. Hydrodynamic studies focus on reductions in drag, while physiological research conclude that drafting lowers metabolic costs. Performance predictions could therefore benefit from incorporating passive drag measures. External conditions, such as wetsuit use (unobservable, if allowed in an event and race format subject to race organizers and regulations), technological wetsuit evolution and innovative race formats (such as the ’Super Sprint’ e.g.) or short-term race format shocks (such as the Covid-19 pandemic), further influence drafting effectiveness. Energy conservation, race strategy, and equipment continues to shape training methodologies and race regulations in the modern triathlon settings.
@@
The dataset originates from three relational database table dumps provided by the Triathlon Statistik Austria database, covering most official triathlon races that took place in Austria from 2010 to 2024.
@@
Table 1: Dataset Description
Variable Description Indexing
events ($\textbf{events}(e,t)$)
event_id Unique event identifier $e$ (event)
date Date of event $t$ (time)
event_category Event type (Sprint, Kurz, etc.) $e$ (event)
location Event location $e$ (event)
athlete ($\textbf{athlete}(a)$)
athlete_id Unique athlete identifier $a$ (individual)
year Year of birth $a$ (individual)
male Binary indicator (1 = Male, 0 = Female) $a$ (individual)
result ($\textbf{result}(r,a,e,t)$)
athlete_id Unique athlete identifier $a$ (individual)
event_id Event identifier $e$ (event)
$totalTime_{a,r,e,t}$ Total race time (seconds) $a,r,e,t$
$swimOutTime_{a,r,e,t}$ Swim segment time (seconds) $S_{a,r,e,t}$
$raceRank_{a,r,e,t}$ Athlete’s rank in event $a,r,e,t$
@@
  • Merging Data: Joined result with athlete and events using athlete_id and event_id.
  • Handling Missing Values: Dropped rows with missing values.
  • Period Assignment: Events were assigned to one of three periods based on their date:
    $P_{\text{pre\_covid}}$ =$=\{t\mid\text{year}_{e}(t)<2020\}$
    $P_{\text{covid}}$ =$=\{t\mid 2020\leq\text{year}_{e}(t)\leq 2022\}$
    $P_{\text{post\_covid}}$ =$=\{t\mid\text{year}_{e}(t)>2022\}$
  • Computing Age (Encoded numerical example variable — used as distributional covariate balance check across periods):
    $\text{age}_{a,t}=\text{year}_{e}(t)-\text{year}_{a}$
@@
To analyze drafting effects, swimmers were clustered based on their swim-out time within an event. A hierarchical clustering algorithm was applied with a 5-second threshold for swim group formation.
@@
The applied agglomerative hierarchical clustering procedure can be formalised as follows:
For each event $e$, let $S_{a,r,e,t}$ be the swim-out time for athlete $a$ in result $r$. The clustering procedure follows these steps:
1. Compute the distance matrix $D$ using Euclidean distance:
$D(a,a^{\prime})=|S_{a,r,e,t}-S_{a^{\prime},r,e,t}|$
for all athletes $a,a^{\prime}$ in event $e$.
2. Perform hierarchical clustering using the cluster single linkage method:
$\delta(A,B)=\min\{D(a,b)\mid a\in A,b\in B\}$
3. Apply a cutoff threshold of 5 seconds:
$C_{a,t}=\begin{cases}c,&\text{if }|S_{a,r,e,t}-S_{a^{\prime},r,e,t}|<5\text{ sec}\\ c^{\prime},&\text{otherwise}\end{cases}$
where athletes within 5 seconds are assigned the same cluster (swim group).
4. Assign a leader and drafters:
$\text{leader}_{a,t}=\begin{cases}1,&\text{if }S_{a,r,e,t}=\min(S_{a^{\prime},r,e,t})\text{ in cluster}\\ 0,&\text{otherwise}\end{cases}$
$\text{drafter}_{a,t}=1-\text{leader}_{a,t}$
5. Compute individual drafting positions within each cluster (swim group):
[…]
@@ 5. Compute individual drafting positions within each cluster (swim group):
Balance checks across event categories and periods
@@
Event Category Period Mean Swim-Out Time SD Swim-Out Time Count
@@
Figure 1: Swim Out Times by Event Category and Period Balance Check
Balance check of athletes across periods
Table 4: Athletes across periods
Period Mean Total Mean Swim-Out Time Mean Rank Mean Age Mean Male Count
Covid 12091 1705 121 40.2 0.768 18467
Post-Covid 12573 1765 127 40.8 0.757 14419
Pre-Covid 11665 1599 165 38.1 0.796 135505
Figure 2: Age Distribution by Period
@@
This study leverages the staggered race starts implemented during the COVID-19 pandemic to estimate the magnitude of open-water swim drafting benefits. Before 2020, athletes started in groups, allowing drafting effects to influence race performance. However, during the pandemic, individually staggered race starts significantly reduced drafting opportunities. This natural experiment provides a unique opportunity to assess the impact of drafting by comparing race outcomes before, during, and after the pandemic.
To formally identify the effects of drafting, I estimate the impact of swim group position on Swim-Out times, focusing on clustered swim-out times. I furthermore employ athlete and event fixed effects to control for unobservable heterogeneity and to address some potential endogeneity and reverse causality concerns. The identification strategy, scenarios, and edge cases are illustrated in subsequent subsections.
Swim Course and Drafting Positions
The figure below illustrates a hypothetical swim course, where a group of five athletes is clustered based on swim-out times. The numbers indicate the position of each athlete within the group, with trailing athletes experiencing greater drafting benefits.
12345Swim DirectionLeader (No Drafting Benefit)3rd Position (Optimal Drafting)5th Position (High Drafting Benefit)
Figure 3: Example of Swim Group and Drafting Positions
This diagram demonstrates how drafting effects vary based on athlete position within a swim group. The leader experiences no drafting benefits, while trailing athletes in positions 3 to 5 gain progressively more hydrodynamic advantages. The measurements of swim-out-times are only recorded at the end of the swim-segment before proceeding with the bike leg.
Additional Last Drafter Binary Dummy
The sixth swimmer (gray) represents the last drafter, still potentially gaining higher hydrodynamic benefits while being assigned to the same cluster (swim group).
123456Swim Direction
Figure 4: Swim Group with a last drafter - in this example also the fifth drafter.
Triangular Swim Course
In a triangular swim course is very typical in triathlon races, athletes navigate around two buoys in a triangular path. The buoys create tactical points where overtaking is challenging, emphasizing positioning to avoid potential hits and washing machine drowning dynamics. Sometimes multiple laps before the swim-out time taking point is reached.
StartBuoy 1Buoy 2
Figure 5: Triangular Swim Course with Buoy Positions
Circular (Round) Swim Course
A circular course is used for many laps-based open-water swimming events. Sometimes multiple laps before the swim-out time taking point is reached.
Figure 6: Circular Swim Course for Open Water Races
Catching Up Scenario
This scenario shows an athlete attempting to catch up to a drafting group. The orange swimmer (C) is closing the gap, potentially affecting positioning. The race dynamics during the swim segment cannot be measured.
123456C
Figure 7: Catching Up Scenario in a Swim Group
No Overtaking at Buoys (COVID-19 Rule in 2020)
Due to COVID-19 restrictions, overtaking at buoys was disallowed in some races to maintain ’social distancing’. The dashed red line marks the restricted overtaking zone.
No Overtaking Zone
Figure 8: No Overtaking at Buoys (2020 COVID-19 Rule)
@@
To estimate the impact of swim drafting on Swim-Out performance, I employ a Two-Way Fixed Effects (TWFE) model, controlling for both athlete- and event-specific heterogeneity. The general estimation equation can be written in the of:
@@
where $y_{aert}$ represents the Swim-Out performance (e.g., swim time) of athlete $a$ in event $e$, race result $r$, at time $t$. The variables $\text{Leader}_{aert}$ and $\text{Drafter}_{aert}^{(p)}$ indicate drafting positions, while the vector $\mathbf{X}_{aert}$ contains relevant controls, including RaceRank or Cluster (Swim Group).
The fixed effects $\eta_{a}$ account for athlete heterogeneity, while $\eta_{et}$ captures event-time-specific shocks such as weather conditions, water temperature, or changes in drafting regulations.
To explore heterogeneity in drafting effects, I extend our baseline specification with interaction terms:
$y_{aert}=\beta_{0}+\beta_{1}\text{Drafter}_{aert}\times\text{EventYear}_{t}+\boldsymbol{\theta}^{\top}\mathbf{X}_{aert}+\eta_{et}+\eta_{a}+\varepsilon_{aert}.$
This model allows us to estimate whether the effect of drafting changed over time, particularly before, during, and after COVID-19, when race conditions varied significantly.
Clustered, Robust and TWFE Standard Errors
To ensure valid statistical inference, I implement a range of robust standard error estimators, accounting for potential heteroskedasticity and within-athlete correlation. The standard error formulations can be written as followed:
@@
To account for heteroskedasticity, I use the HC1 robust variance-covariance estimator:
$\hat{\Sigma}_{HC1}=\left(\sum_{a,e,t}\mathbf{X}_{aert}^{\top}\mathbf{X}_{aert}\right)^{-1}\left(\sum_{a,e,t}\mathbf{X}_{aert}^{\top}\hat{\varepsilon}_{aert}^{2}\mathbf{X}_{aert}\right)\left(\sum_{a,e,t}\mathbf{X}_{aert}^{\top}\mathbf{X}_{aert}\right)^{-1}$
@@
To account for within-athlete correlation across multiple events, I cluster standard errors at the athlete level:
$\hat{\Sigma}_{\text{Athlete}}=\left(\sum_{a,e,t}\mathbf{X}_{aert}^{\top}\mathbf{X}_{aert}\right)^{-1}\left(\sum_{a}\mathbf{X}_{aert}^{\top}\hat{\varepsilon}_{aert}\hat{\varepsilon}_{aert}^{\top}\mathbf{X}_{aert}\right)\left(\sum_{a,e,t}\mathbf{X}_{aert}^{\top}\mathbf{X}_{aert}\right)^{-1}$
This accounts for athlete-specific dependencies and ensures correct inference when an athlete competes in multiple events.
@@
To address potential correlation within both athlete and event clusters, I use multi-way clustering:
$\hat{\Sigma}_{\text{Athlete, Event}}=\hat{\Sigma}_{\text{Athlete}}+\hat{\Sigma}_{\text{Event}}-\hat{\Sigma}_{\text{Athlete}\cap\text{Event}},$
where:
$\hat{\Sigma}_{\text{Event}}=\left(\sum_{e}\mathbf{X}_{aert}^{\top}\hat{\varepsilon}_{aert}\hat{\varepsilon}_{aert}^{\top}\mathbf{X}_{aert}\right)$
This accounts for common shocks affecting all athletes in a given event, such as weather conditions, drafting regulations, drafting group dynamics, and water temperature.
@@
To further account for unobserved heterogeneity across both athlete and event dimensions, I employ the Two-Way Fixed Effects (TWFE) variance estimator:
$\hat{\Sigma}_{TWFE}=\left(\sum_{a,e,t}\mathbf{X}_{aert}^{\top}M_{\eta}\mathbf{X}_{aert}\right)^{-1}\left(\sum_{a,e,t}\mathbf{X}_{aert}^{\top}M_{\eta}\hat{\varepsilon}_{aert}^{2}M_{\eta}\mathbf{X}_{aert}\right)\left(\sum_{a,e,t}\mathbf{X}_{aert}^{\top}M_{\eta}\mathbf{X}_{aert}\right)^{-1}$
where $M_{\eta}$ is the demeaning operator that removes athlete and event fixed effects.
By employing these various robust standard error estimators, our analysis ensures statistically valid inference, reducing the risk of misleading standard errors due to intra-cluster correlations.
Empirical Findings
Findings I.
Individually staggered race starts during the pandemic and precautionary COVID-19 measures rendered the benefits of open-water swim drafting statistically insignificant in 2020. In the years following the pandemic, the estimated effects of drafting appear to be somewhat less pronounced. This result holds particularly for a subsample consisting of athletes who competed across all time periods. (See Appendix)

### v2

# Using Covid-19 Response Policy to Estimate Open Water and Swim Drafting Effects in Triathlon

@@ Abstract
Objectives
This study investigates the causal effects of open-water swim drafting by leveraging a natural experiment induced by staggered race starts during the COVID-19 pandemic. Furthermore, the study aims to quantify the benefits of swim drafting in real-world triathlon race settings and estimate the effect of different drafting positions within swim groups.
Design
A quasi-experimental design is employed, exploiting exogenous variation in swim drafting opportunities caused by COVID-19 restrictions. The study uses panel data from triathlon races before, during, and after the pandemic.
Methods
Using agglomerative hierarchical clustering, likely swim group formations are computed from swim-out segment times. The impact of drafting positions on swim performance is estimated using a two-way fixed-effects regression model, controlling for athlete- and event-specific factors. Robustness checks in the form of alternative regression model specifications and samples are used to partially address some potential endogeneity concerns, including reverse causality and omitted variable bias.
Results
Empirical findings reveal that drafting benefits were statistically insignificant in 2020 due to reduced drafting opportunities but re-emerged post-pandemic at slightly lower levels than pre-pandemic periods. The results indicate that drafting only becomes beneficial from the third trailing position onward, with earlier positions primarily serving to minimize fatigue, as shown by an observed inverse decay in drafting benefits (higher estimated penalties in lower-order drafting positions/stronger higher-order drafting effects).
Conclusions
This study provides the first large-scale causal estimate of open-water swim drafting effects in real triathlon race conditions. The findings contribute to the understanding of race strategy optimization (metabolic costs), offering new insights for endurance athletes, race organizers and regulation authorities. Additionally, the results indicate the importance of strategic positioning within swim groups to maximize performance benefits while mitigating fatigue effects.
@@ Introduction
In these latter formats, the only real opportunity for drafting occurs in the swim segment, where swimmers can strategically position themselves to reduce drag. Swim drag reduction through drafting has been extensively studied in medical science, soft matter physics and sports science in Bolon et al. (2020); Chatard et al. (1990); Chatard and Wilson (2003); Ohmichi et al. (1983); Millet et al. (1999); Delextrat et al. (2003); Olbrecht (2011); Toussaint et al. (1989). However, no large-scale causal estimates of its benefits in real race settings have been published.
The COVID-19 pandemic introduced precautionary measures in Austria which significantly altered the swim segment dynamics. According to Österreichischer Triathlonverband (ÖTRV) Athletes often had to start with individually staggered swim entries, and overtaking at buoy turns was sometimes somewhat restricted. This created a natural experiment where COVID-19 functioned as an exogenous short-term shock (1–3 years) to open-water swim drafting, providing an opportunity to analyze its effects in a controlled manner.
@@
Water drafting has long been a hot topic in both sports science and professional triathlon. With internet sources claiming that an athlete can gain up to between $\approx$ 7%-21% on overall drag reduction, which results in a reduction of 2 minutes over a half-distance Ironman swim of 1,900m. [^1] Drafting in Triathlon: https://support.myprocoach.net/hc/en-us/articles/360022528551-Drafting-in-Triathlon#:~%3Atext=Swim%20Drafting,you%20up%20to%20two%20minutes.
Beyond race performance, other strategic considerations—such as energy conservation and positioning to catch the first cycling group in elite competitions where wind drafting is permitted—play a critical role.
@@
Hydrodynamic and Physiological Benefits of Drafting. Studies have shown that drafting significantly reduces drag and physiological strain. Bolon et al. (2020) determined that optimal drafting positions—directly behind a lead swimmer or at the hip—result in drag reductions of 40% and 30%, respectively. Chatard and Wilson (2003) found that metabolic costs, including oxygen uptake, heart rate, and blood lactate levels, were significantly lower when swimmers drafted at distances of 0 to 50 cm behind another swimmer’s toes. The most advantageous drafting distances were 0 and 50 cm, reducing passive drag by 21% and 20%, respectively. Additionally, Ohmichi et al. (1983) found that swimmers drafting directly behind another competitor completed 400-m races significantly faster than those who swam at the side or hip.
@@
Drafting and Performance Prediction. The role of passive drag (Dp) in swim performance prediction was examined by Chatard et al. (1990), who found a strong correlation between Dp and VO2 max. Performance times were primarily influenced by VO2 max ($r=0.70$ for males, $r=0.72$ for females), but including Dp significantly improved the predictive accuracy. The strong correlation hints at the possibility that more able athletes might be more sensitive to properly utilizing passive drag. Other anthropometric factors, including height, weight, and body surface area, also contributed to drag variability. This athlete heterogeneity underscores the need to include athlete-specific intercepts or characteristics in predictive models. Millet et al. (1999) further emphasized the cumulative impact of drafting, showing that triathletes employ energy-conserving strategies in swimming to enhance their subsequent cycling and running performances.
@@
Tactical Implications for Triathletes. Drafting has been shown to enhance subsequent cycling performance. Delextrat et al. (2003) reported a 4.8% improvement in cycling efficiency when a swim was performed in a drafting position compared to an isolated swim. Millet et al. (1999) also stated the cumulative effects of drafting in triathlon, where athletes strategically conserve energy in the swim segment to optimize performance in the cycle and run segment. Olbrecht (2011) thus underscores the importance of balancing conditioning and technique, arguing that poor swimming technique results in excessive energy expenditure, ultimately affecting overall race outcomes.
@@
Impact of Equipment and External Conditions. The effect of wetsuits on swim drafting was investigated by Toussaint et al. (1989), who observed a 14% reduction in drag at typical triathlon swim speeds. This reduction, largely attributed to increased buoyancy and decreased frontal resistance, resulted in higher swim velocities. Moreover, Leitner (2021) conducted an athlete survey revealing preferences regarding mass starts versus staggered starts, highlighting the impact of the race format on drafting opportunities. An alternative survey on wetsuit use could further reveal preferences for drafting but is likely highly confounded by the swim-to-bike transitional race segment. The COVID-19 pandemic introduced natural experiments with staggered starts, as documented by Österreichischer Triathlonverband (ÖTRV), which temporarily altered swim drafting dynamics.
@@
Summary of Existing Literature. The body of research highlights drafting as a critical factor in competitive swimming and triathlon. Hydrodynamic studies focus on reductions in drag, while physiological research conclude that drafting lowers metabolic costs. Performance predictions could therefore benefit from incorporating passive drag measures. External conditions, such as wetsuit use (unobservable, if allowed in an event and race format subject to race organizers and race regulation authorities), technological wetsuit evolution and innovative race formats (such as the ’Super Sprint’ e.g.) or short-term race format shocks (such as the Covid-19 pandemic), further influence drafting effectiveness.
@@
The dataset originates from three relational database tables provided by the Triathlon Statistics Austria database, covering most official triathlon races that took place in Austria from 2010 to 2024.
@@
The dataset was processed by first merging the result table with the events and athlete table using foreign keys. Events were then assigned to one of three periods based on their date, resulting in three binary dummy variables. Additionally, event-time-variant covariates such as age or squared age were computed in the final dataset as $\text{age}_{a,t}:=\text{year}_{e}(t)-\text{year}_{a}$, which were then additionally used as a distributional covariate balance check across periods.
@@
To analyze drafting effects, swimmers were clustered based on their swim-out time within an event. A hierarchical clustering algorithm was applied with a 5-second threshold for forming swim groups.
@@
The clustering procedure involved calculating the time difference between athletes in an event and grouping them based on proximity. A single linkage method was used, ensuring that swimmers within 5 seconds of each other were assigned to the same swim group.
Each cluster was assigned a leader, defined as the athlete with the fastest swim-out time. All other athletes in the group were designated as drafters. Within the drafting group, individual positions were encoded, categorizing athletes as first, second, third, fourth, fifth, and/or last drafters based on their relative swim-out times.
@@
Balance Checks
To ensure comparability (e.g. problems with the panel data such as attrition) across event categories and periods, some descriptive and visual balance checks on key event, result and athlete variables were performed.
Swim-out times across Event Category and Period Balance
Swim-out times were analyzed across different event categories (Sprint, Short, Middle, and Long) and periods (Pre-Covid, Covid, and Post-Covid) to assess differences. The results indicate slight increases in mean swim-out times post-Covid for most Sprint and Short events. This trend however was expected, as the movement towards more accurate swim distances and longer swim distances in the Short category over the last decade is well known among Austrian triathlon insiders. The balance check is shown in table 2.
Athlete Characteristics Across Periods
Athlete-level characteristics, including mean swim-out time, rank, age, and the proportion of male participants, were examined to assess balance across periods. Age distribution trends indicate a slight increase in the average age of athletes since the pre-Covid period starting in 2010, alongside a slight decline in the overall high proportion of male participants. Additionally, the observed decrease in mean rank suggests a shift in event market structure, potentially favoring a higher number of smaller events or fewer mass events. However, this trend could also be entirely driven by an outlier event within the longer period spanning 2010 to 2019.
@@
Event Category Period Mean Swim-Out Time SD Swim-Out Time Observations
@@
Event-Induced Altered Drafting Opportunity.
This study leverages the staggered race starts implemented during the COVID-19 pandemic to estimate the magnitude of drafting benefits in open-water swimming. Before 2020, athletes started in groups, allowing drafting effects to influence race performance. However, during the pandemic, individually staggered starts significantly reduced drafting opportunities. This natural experiment offers a unique opportunity to assess drafting’s impact by comparing race outcomes before, during, and after the pandemic.
Confoundedness, Effect Heterogeneity and Higher-Order Drafting Position Effects.
To formally identify the effects of drafting, I estimate the impact of drafting and within-swim-group position on Swim-Out times, using likely constructed swim groups based on clustered Swim-Out times. Furthermore, I employ athlete and event fixed effects to control for unobservable heterogeneity and to partially address some endogeneity concerns such as omitted variable bias and simultaneous reverse causality.
The positioning within a swim group likely affects the magnitude of drafting benefits, with higher-order positions benefiting more from overall reduced water resistance (increased passive drag), leading to varying performance outcomes. However, reverse causality is possible if swimmers in the last positions within swim groups only manage to catch up towards the end of the swim segment and are, in general, more skilled and/or performing athletes. Additionally, a strong correlation between passive drag and $\dot{V}O_{2\max}$ could also indicate that stronger athletes also experience greater passive drag Chatard et al. (1990), assuming this result applies also to real-world race settings.
Unobservables and COVID-19 Event Regulations.
Single catching-up scenarios cannot be fully accounted for in the estimated models. The potential bias introduced by such instances is assumed to be uniformly distributed across all race observations.
Special consideration is given to regulatory constraints, such as the COVID-19 rule prohibiting overtaking in specific course segments. These restrictions further influence race drafting opportunities throughout the swim segment, ultimately affecting overall mean race outcomes for drafters. Additionally, different observable course configurations, such as triangular and circular swim courses, shape drafting benefits and race dynamics.
Furthermore, Buoy placements create tactical points that influence overtaking and positioning strategies, particularly in multi-lap races where swimmers navigate the same course multiple times before exiting the swim segment.
@@
To estimate the impact of swim drafting on Swim-Out performance, I employ a Two-Way Fixed Effects (TWFE) model, controlling for both athlete- and event-specific heterogeneity. The general estimation equation can be written in the form of:
@@
where $y_{aert}$ represents the Swim-Out performance (e.g., swim time) of athlete $a$ in event $e$, stored as race result $r$, at time $t$. The encoded binary dummy variables $\text{Leader}_{aert}$ and $\text{Drafter}_{aert}^{(p)}$ indicate drafting positions, while the vector $\mathbf{X}_{aert}$ contains relevant controls, including the overall race rank or constructed cluster rank (Swim Group) in the swim segment.
The fixed effects $\eta_{a}$ account for athlete heterogeneity, while $\eta_{et}$ captures event-time-specific shocks such as weather conditions, water temperature, or changes in race regulations. (e.g. wet suit allowance)
To explore period heterogeneity in drafting effects, I estimate a second specification with interaction terms:
$y_{aert}=\beta_{0}+\text{Drafter}_{aert}\times(\beta_{1}\text{PrePeriod}_{t}+\beta_{2}\text{CovidYear}_{t})+\boldsymbol{\theta}^{\top}\mathbf{X}_{aert}+\eta_{et}+\eta_{a}+\varepsilon_{aert}$
This model allows us to estimate whether the effect of drafting changed over time, particularly before, during, and after COVID-19, when race regulations varied significantly.
Clustered, Robust, and TWFE Standard Errors
To ensure reliable statistical inference, I employ various robust standard error estimators that account for heteroskedasticity and correlations within athletes across multiple events.
@@
To address potential heteroskedasticity—where the variability of errors differs across observations—I use a robust standard error estimator. This approach adjusts the standard errors to remain valid even when the underlying variance structure is inconsistent across data points.
@@
To account for the fact that individual athletes may participate in multiple events, I cluster standard errors at the athlete level. This method ensures that the estimated errors account for within-athlete correlations, leading to more accurate inference when analyzing repeated observations from the same athlete.
@@
Since dependencies may arise not only within athletes but also within events, I implement a multi-way clustering approach. This accounts for common influences affecting all participants in a specific event, such as environmental factors, drafting regulations, group dynamics, and other shared conditions. By doing so, the standard error estimates better reflect the true underlying variation present in the data.
@@
To further control for unobserved heterogeneity across both athlete and event dimensions, I employ the Two-Way Fixed Effects (TWFE) variance estimator. This approach removes systematic differences associated with individual athletes and specific events, ensuring that standard error estimates are robust to persistent unobserved factors. By eliminating these fixed effects, the estimator better isolates the true effect of interest while maintaining statistical validity.
Results
The Covid-19 Response Policy Induced Event
During the pandemic, individually staggered race starts and precautionary COVID-19 measures rendered the benefits of open-water swim drafting statistically insignificant in 2020. In the years following the pandemic, the estimated effects of drafting appear to be somewhat less pronounced. This trend is found to be particularly evident in a subsample of athletes who competed across all time periods, potentially due to COVID-19-related health shocks, changes in training behavior, or panel data limitations such as attrition. (See Appendix)
Figure 1: Drafting benefits based on non-random yearly sub-samples: Coefficient Plot. The solid thin line at 0 represents the baseline with no effect. The dashed vertical line at $x=2020$ marks the year when mass starts were disallowed at many events as a precautionary measure due to the COVID-19 pandemic. Furthermore COVID-19 related race regulations were in place. In this year the coefficient of computed drafter is found to be insignificant.